OpenStudy (anonymous):

choose h and k such that the system has (a) no solutions, (b) a unique solution and (c) many solutions. 1 -3 1 2 h k

myininaya (myininaya):

Tech your question seems like it is lacking something. I can't do anything with this :(

OpenStudy (anonymous):

i made it to matrix form

OpenStudy (anonymous):

original equations are x1-3x2=1 2x1+hx2=k

myininaya (myininaya):

Oh that is a matrix?

OpenStudy (anonymous):

yes

myininaya (myininaya):

ok. \[x-3y=1\] \[2x+hy=k\] ---------------------------- if we want no solution we want the lines to be parallel. So we want the same slope. \[y=\frac{1}{3}x-\frac{1}{3}\] \[y=\frac{-2x}{h}+\frac{k}{h}\] So we want 1/3=-2/h <-------solve for h and this will give you the h so that the two lines that never cross. ----------------------------- now we want there to be infinity many solutions which means we want the lines to lay on top of each other so we want same slope and same y-intercept so we have 1/3=-2/h and -1/3=k/h <---solve these two equations -----------------------------

myininaya (myininaya):

now for one solution that means we want different slope with any y-intercept so we want \[\frac{1}{3} \neq \frac{-2}{h}\]